# Quantitative estimates for singularity for conjugate equations driven by linear fractional transformations

### v3

@@ 3. Singularity
In this section, we consider (1.1) under the setting that $\{g_{i}\}_{i}$ is an LF-system on $[0,1]$, and, $\{f_{i}\}_{i}$ is a weakly contractive D-system on $[0,1]$. By Lemma 2.8, both of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are D-systems. Hence, by Lemma 2.3, there exists a unique continuous strictly increasing solution $\varphi$ of (1.1).
We consider singularity for the solution $\varphi$. We will give sufficient conditions for $\dim_{H}\mu_{\varphi}<1$. It is known that if $\dim_{H}\mu_{\varphi}<1$, then, $\varphi$ is a singular function on $[0,1]$, specifically, $\varphi$ is a strictly increasing and continuous function on $[0,1]$ such that its derivative is equal to zero almost everywhere.
The main purpose of this section is to extend [27, Theorem 43] to the case that $f_{i}$ can be non-linear. In [27], the author dealt with many objects other than the conjugate functional equation, so the arguments were not specialized for the functional equation. We introduce some new notations and the formulations of our main results and some of the arguments in the proofs are different from [27, Theorem 43]. These changes will make arguments more transparent.
We assume that an LF-system $\{g_{i}\}_{i}$ is defined by $g_{i}(x):=\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1]$, $i\in\mathcal{I}_{N}$, where $a_{i},b_{i},c_{i},d_{i},i\in\mathcal{I}_{N}$, are real numbers satisfying the conditions (1), (2) and (3) in Definition 2.7. In this section, we assume $d_{i}=1$ for each $i\in\mathcal{I}_{N}$.
@@
$\alpha:=\min\left\{0,\frac{c_{0}}{1-a_{0}},\frac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\middle|1\leq i\leq N-1\right\},$
where we delete [^2] This is different from [27]. $\frac{c_{0}}{1-a_{0}}$ if $a_{0}=1$. Let
$\beta:=\max\left\{0,\frac{c_{0}}{1-a_{0}},\frac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\middle|1\leq i\leq N-1\right\},$
where $\dfrac{c_{0}}{1-a_{0}}:=+\infty$ if $a_{0}=1$.
Let $Y:=[\alpha,\beta]$. Let
$G_{i}(y):=\frac{(a_{i}-b_{i}c_{i})(y+1)}{(b_{i}y+1)((a_{i}+b_{i})y+(c_{i}+1))}=\frac{(b_{i+1}-b_{i})(y+1)}{(b_{i+1}y+1)(b_{i}y+1)},\ y\in[\alpha,\beta].$
@@
$H_{i}(y):=\frac{a_{i}y+c_{i}}{b_{i}y+1},\ y\in[\alpha,\beta].$
Let $G_{i}(+\infty)$ and $H_{i}(+\infty)$ be the limits. Specifically,
$G_{0}(+\infty):=1,G_{i}(+\infty):=0,H_{0}(+\infty):=+\infty,H_{i}(+\infty):=\frac{a_{i}}{b_{i}},1\leq i\leq N-1.$
@@
(iv) $\sum_{i\in\Sigma_{N}}G_{i}(y)=1,\ y\in Y.$
By (iii), we can compose $H_{i}$ freely on $Y$. We consider the compactification when $\beta=+\infty$. Then, $Y=[\alpha,\beta]$ is compact, and $G_{i}:Y\to[0,1]$ and $H_{i}:Y\to Y$ are continuous. We remark that $Y,G_{i},H_{i}$ are defined by $(g_{i})_{i}$, not by $(f_{i})_{i}$.
$\textup{Fix}(H_{i})$ denotes the set of fixed points of $H_{i}$ on $[\alpha,\beta]$. This is uniquely determined and
@@
$\mathcal{P}_{N}:=\left\{(p_{i})_{i\in\mathcal{I}_{N}}:p_{i}>0,\sum_{i}p_{i}=1\right\}$
and
$\overline{\mathcal{P}_{N}}:=\left\{(p_{i})_{i\in\mathcal{I}_{N}}:p_{i}\geq 0,\sum_{i}p_{i}=1\right\}.$
@@
${\bf G}(y):=(G_{0}(y),\dots,G_{N-1}(y))\in\overline{\mathcal{P}_{N}},\ y\in Y.$
Let
$\mathfrak{p}_{i}:={\bf G}\left(\textup{Fix}(H_{i})\right)\in\overline{\mathcal{P}_{N}},\ i\in\mathcal{I}_{N}.$
Let
$\dim_{H}\mu_{\varphi}:=\inf\left\{\dim_{H}A:A\in\mathcal{B}([0,1]),\mu_{\varphi}(A)=1\right\}.$
The following results extend [27, Theorem 43 (i)] to the case that $\{f_{i}\}_{i}$ may be non-affine.
@@
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|\geq 2$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i\in\mathcal{I}_{N}}r_{i}<1+\delta$ and $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$, where we let $r_{i}:=\textup{Lip}(f_{i})$.
The probability vectors $\mathfrak{p}_{i},i\in\mathcal{I}_{N}$, are determined only by $(\alpha,\beta,\{G_{i}\}_{i},\{H_{i}\}i)$ and date back to $\{g_{i}\}_{i}$. $\{f_{i}\}_{i}$ is not related to this. This contains the affine case that $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i\in\mathcal{I}_{N}$. The reason why we have introduced $Y,G_{i}$ and $H_{i}$ will be clarified in the proof.
We also deal with the case that |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$. Let the total distance between probability vectors be
$\|{\bf p}-{\bf q}\|:=\sum_{i\in\mathcal{I}_{N}}|p_{i}-q_{i}|,\ \ {\bf p}=(p_{i})_{i},{\bf q}=(q_{i})_{i}\in\overline{\mathcal{P}_{N}}.$
@@
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i\in\mathcal{I}_{N}}r_{i}<1+\delta$, $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$, and $\|{\bf r}-\mathfrak{p}_{0}\|\geq\epsilon$, where we let $r_{i}:=\textup{Lip}(f_{i})$ and ${\bf r}:=(r_{0},\cdots,r_{N-1})$.
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, $\mathfrak{p}_{0}=\cdots=\mathfrak{p}_{N-1}$. If $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i\in\mathcal{I}_{N}$, and $\mathfrak{p}_{0}={\bf q}$, then, the solution $\varphi$ is smooth on $[0,1]$ and we also have an explicit expression for $\varphi$. See [27, Theorem 43 (ii)].
@@ 3.1. Proofs
We first give notations for symbolic dynamics. Let :=$\mathbb{N}:=\{1,2,\cdots\}$. We define a map $\pi:(\mathcal{I}_{N})^{\mathbb{N}}\to[0,1]$ by
$\pi(i_{1}i_{2}\cdots):=\lim_{k\to\infty}f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0).$
This is well-defined because each $f_{i}$ is a contraction and $\left(f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0)\right)_{k}$ is a Cauchy sequence.
Let $\xi_{j}:(\mathcal{I}_{N})^{\mathbb{N}}\to\mathcal{I}_{N}$ be the projection to the $j$-th coordinate. Let
[…]
@@
$\pi\left(I(i_{1},\cdots,i_{k})\right)=\left[f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0),f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(1)\right].$
@@
$\mu_{\varphi}\left(\pi(I(i_{1},\cdots,i_{k}))\right)$ $=\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{k}}(1)\right)-\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{k}}(0)\right)$
$=g_{i_{1}}\circ\cdots\circ g_{i_{k}}(1)-g_{i_{1}}\circ\cdots\circ g_{i_{k}}(0).$
@@
$\frac{\nu_{y}(I(i_{1},\cdots,i_{n-1},i_{n}))}{\nu_{y}(I(i_{1},\cdots,i_{n-1}))}=G_{i_{n}}\circ H_{i_{n-1}}\circ\cdots\circ H_{i_{1}}(y),\ n\geq 2.$ (3.1)
Since $\sum_{i\in\mathcal{I}_{N}}G_{i}(y)=1$ for each $y\in Y$, the consistency condition holds. Therefore, a probability measure on $(\mathcal{I}_{N})^{\mathbb{N}}$ satisfying (3.1) uniquely exists by the Kolmogorov extension theorem.
For $i\in\mathcal{I}_{N}$, let $F_{i}:(\mathcal{I}_{N})^{\mathbb{N}}\to(\mathcal{I}_{N})^{\mathbb{N}}$ be a map defined by $F_{i}(z)=iz$. Now we see that
$\nu_{y}=\sum_{i\in\mathcal{I}_{N}}G_{i}(y)\nu_{H_{i}(y)}\circ F_{i}^{-1}.$ (3.2)
Let $\mu_{y}:=\nu_{y}\circ\pi^{-1}$. Then, $\mu_{y}$ is a Borel probability measure on $[0,1]$. By $\pi\circ F_{i}=f_{i}\circ\pi$ and (3.2),
$\mu_{y}=\sum_{i\in\mathcal{I}_{N}}G_{i}(y)\mu_{H_{i}(y)}\circ f_{i}^{-1}.$ (3.3)
@@
Since $H_{i}(0)\neq 0$ can occur, this does not define an iterated function system. (3.3) is in the framework of a place-dependent iterated function system [9].
@@
$\nu_{0}\circ\pi^{-1}=\mu_{0}=\mu_{\varphi}.$ (3.4)
Let $\epsilon\in(0,1/2)$. We assume that $\epsilon\leq\min_{i\in\mathcal{I}_{N}}r_{i}\leq\max_{i\in\mathcal{I}_{N}}r_{i}\leq 1-\epsilon$. Since $\{f_{i}\}_{i}$ is a compatible system, $\sum_{i\in\mathcal{I}_{N}}r_{i}\geq 1$ and $r_{i}>0$ for each $i\in\mathcal{I}_{N}$. Let $s\geq 1$ is a constant such that $\sum_{i\in\mathcal{I}_{N}}r_{i}^{s}=1$.
@@
Let $\epsilon_{2}>0$. Then, for $\delta>0$ such that
$\frac{\log N}{\log(N/(1+\delta))}<1+\frac{\epsilon_{2}}{\log 1/\epsilon},$ (3.5)
it holds that $s<1+\dfrac{\epsilon_{2}}{\log 1/\epsilon}$ for every choice of $(r_{i})_{i\in\mathcal{I}_{N}}\in\mathcal{P}_{N}$ such that $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$.
@@
Let $u_{i}:=r_{i}/\sum_{i}r_{i}$. Then, $\sum_{i}u_{i}=1$. We see that $\sum_{i}r_{i}^{s}=1$ if and only if $\sum_{i}u_{i}^{s}=(\sum_{i}r_{i})^{-s}$. Hence, if $\sum_{i}r_{i}\leq 1+\delta$, then,
$\frac{\log(\sum_{i}u_{i}^{s})}{s}\leq-\log(1+\delta).$
$\sum_{i}u_{i}^{s}$ takes its minimum at $u_{0}=\cdots=u_{N-1}=\dfrac{1}{N}$. We see this by the method of Lagrange multipliers. The minimum is $N^{1-s}\leq 1$. Hence, $-\dfrac{s-1}{s}\log N\geq-\log(\sum_{i}r_{i})$, which is equivalent to $s<\dfrac{\log N}{\log(N/(1+\delta))}$. Therefore, if (3.5) holds, then, $s<1+\dfrac{\epsilon_{2}}{\log 1/\epsilon}$. ∎
@@
$R_{n}({\bf q};z):=\frac{\nu_{0}(I_{n}(z))}{q_{\xi_{1}(z)}\cdots q_{\xi_{n}(z)}},\ \ {\bf q}=(q_{i})_{i}\in\mathcal{P}_{N},z\in(\mathcal{I}_{N})^{\mathbb{N}}.$
By a measure-theoretic argument, we can show that
@@
Let $\widetilde{q}_{i}:=r_{i}^{s}>0$ and ${\bf\widetilde{q}}:=(\widetilde{q}_{i})_{i}\in\mathcal{P}_{N}$. If there exists $\epsilon_{2}>0$ such that
$\limsup_{n\to\infty}\frac{-\log R_{n}({\bf\widetilde{q}};z)}{n}\leq-\epsilon_{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$ (3.6)
then,
@@
We omit the proof of this assertion. See the proof of [27, Lemma 7]. (3.4) plays an important role.
By Lemmas 3.4 and 3.5, we obtain that
@@
$\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}\leq-\epsilon_{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$ (3.7)
Then, for $\delta=\delta(N,\epsilon,\epsilon_{2})>0$ satisfying (3.5), $\dim_{H}\mu_{\varphi}<1$.

### v4

@@ 3. Singularity
In this section, we consider (1.1) under the setting that $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ is an LF-system on $[0,1]$, and, $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ is a weakly contractive D-system on $[0,1]$. By Lemma 2.7 and Lemma 2.9, both $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are D-systems. Hence, by Lemma 2.3, there exists a unique continuous strictly increasing solution $\varphi$ of (1.1).
We consider the singularity for the solution $\varphi$. We will give sufficient conditions for $\dim_{H}\mu_{\varphi}<1$. It is known that if $\dim_{H}\mu_{\varphi}<1$, then $\mu_{\varphi}$ is singular with respect to the Lebesgue measure. By [28, Definition 9.2.2 and Theorem 9.2.4], this implies that $\varphi$ is a singular function on $[0,1]$, specifically, $\varphi$ is a strictly increasing and continuous function on $[0,1]$ such that its derivative is equal to zero almost everywhere.
The main purpose of this section is to extend [27, Theorem 43] to the case where $f_{i}$ can be non-linear. In [27], the author dealt with many objects other than the conjugate functional equation, so the arguments were not specialized for the functional equation. We introduce some new notations and the formulations of our main results and some of the arguments in the proofs are different from [27, Theorem 43]. The author believes that these changes will make the arguments more transparent.
We assume that an LF-system $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ is defined by $g_{i}(x)\coloneqq\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1],(i\in\mathcal{I}_{N})$, where $a_{i},b_{i},c_{i},d_{i},(i\in\mathcal{I}_{N})$, are real numbers satisfying the conditions (1), (2) and (3) in Definition 2.8, specifically,
(1) It holds that $c_{i}+d_{i}>0$ and $d_{i}>0$ for $i\in\mathcal{I}_{N}$.
(2) It holds that
$0<a_{i}d_{i}-b_{i}c_{i}\leq\min\{d_{i}^{2},(c_{i}+d_{i})^{2}\},\ i\in\mathcal{I}_{N}.$
(3) It holds that $b_{0}=0$ and $\dfrac{a_{i}+b_{i}}{c_{i}+d_{i}}=\dfrac{b_{i+1}}{d_{i+1}}$ for $i\in\mathcal{I}_{N}$, where we let $\dfrac{b_{N}}{d_{N}}\coloneqq 1$.
In this section, we assume $d_{i}=1$ for each $i\in\mathcal{I}_{N}$.
We show that $(1-a_{i})^{2}+4b_{i}c_{i}\geq 0$. If $c_{i}\geq 0$, then this is true since $b_{i}\geq 0$. If $c_{i}<0$, then by the compatibility condition, $a_{i}+b_{i}\leq c_{i}+1$. Hence, $0\leq b_{i}-c_{i}\leq 1-a_{i}$, and then $(1-a_{i})^{2}\geq(b_{i}-c_{i})^{2}$. Hence, $(1-a_{i})^{2}+4b_{i}c_{i}\geq(b_{i}+c_{i})^{2}\geq 0$.
@@
$\alpha\coloneqq\min\left\{0,\frac{c_{0}}{1-a_{0}},\frac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\,\middle|\,1\leq i\leq N-1\right\},$
where we delete [^2] This is different from [27]. $\dfrac{c_{0}}{1-a_{0}}$ if $a_{0}=1$. Let
$\beta\coloneqq\max\left\{0,\frac{c_{0}}{1-a_{0}},\frac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\,\middle|\,1\leq i\leq N-1\right\},$
where $\dfrac{c_{0}}{1-a_{0}}\coloneqq+\infty$ if $a_{0}=1$.
Let $Y\coloneqq[\alpha,\beta]$ and
$G_{i}(y)\coloneqq\frac{(a_{i}-b_{i}c_{i})(y+1)}{(b_{i}y+1)((a_{i}+b_{i})y+(c_{i}+1))}=\frac{(b_{i+1}-b_{i})(y+1)}{(b_{i+1}y+1)(b_{i}y+1)},\ y\in Y,$
as well as
$H_{i}(y)\coloneqq\frac{a_{i}y+c_{i}}{b_{i}y+1},\ y\in Y,$
where we let $b_{N}\coloneqq 1$. We remark that $G_{0}(y)=\dfrac{b_{1}(y+1)}{b_{1}y+1}$ and $G_{N-1}(y)=\dfrac{1-b_{N-1}}{b_{N-1}y+1}$.
@@
$G_{0}(+\infty)\coloneqq 1,G_{i}(+\infty)\coloneqq 0,H_{0}(+\infty)\coloneqq+\infty,H_{i}(+\infty)\coloneqq\frac{a_{i}}{b_{i}},1\leq i\leq N-1.$
Then,
$G_{i}(+\infty)=\lim_{y\to+\infty}G_{i}(y),\ H_{i}(+\infty)=\lim_{y\to+\infty}H_{i}(y),\ i\in\mathcal{I}_{N}.$
@@
(iv) $\displaystyle\sum_{i\in\mathcal{I}_{N}}G_{i}(y)=1,\ y\in Y.$
By (iii), we can compose $H_{i}$ freely on $Y$. We consider the Alexandroff compactification when $\beta=+\infty$. Then, $Y=[\alpha,\beta]$ is compact, and $G_{i}\colon Y\to[0,1]$ and $H_{i}\colon Y\to Y$ are continuous. We remark that $Y,G_{i},H_{i}$ are defined by $(g_{i})_{i}$, not by $(f_{i})_{i}$.
We denote the set of fixed points of $H_{i}$ on $Y$ by $\textup{Fix}(H_{i})$.
The set $\textup{Fix}(H_{i})$ is a singleton and
@@
If $a_{0}=1$, then $c_{0}>0$ and hence $\textup{Fix}(H_{0})=\{+\infty\}$. If $a_{0}=1$ and $c_{0}=0$, then, by using the assumption that $d_{0}=1$, we obtain $g_{0}(x)=x$ and hence $N=1$. This contradicts the assumption that $N\geq 2$.
@@
$\displaystyle\mathcal{P}_{N}$ $\displaystyle\coloneqq\left\{(p_{i})_{i\in\mathcal{I}_{N}}\colon p_{i}>0,\sum_{i}p_{i}=1\right\}\ \textup{ and}$
$\displaystyle\overline{\mathcal{P}_{N}}$ $\displaystyle\coloneqq\left\{(p_{i})_{i\in\mathcal{I}_{N}}\colon p_{i}\geq 0,\sum_{i}p_{i}=1\right\}.$
@@
$\displaystyle{\bf G}(y)$ $\displaystyle\coloneqq(G_{0}(y),\dots,G_{N-1}(y))\in\overline{\mathcal{P}_{N}},\ y\in Y,$
$\displaystyle\mathfrak{p}_{i}$ $\displaystyle\coloneqq{\bf G}\left(\textup{Fix}(H_{i})\right)\in\overline{\mathcal{P}_{N}},\ \ \ \ \ \ \ \ \ \ \ \ \ \ i\in\mathcal{I}_{N},\ \ \ \ \ \ \ \ \ \textup{ and,}$
$\displaystyle\dim_{H}\mu_{\varphi}$ $\displaystyle\coloneqq\inf\left\{\dim_{H}A\colon A\in\mathcal{B}([0,1]),\mu_{\varphi}(A)=1\right\},$
where $\mathcal{B}([0,1])$ is the Borel $\sigma$-algebra on $[0,1]$. The following results extend [27, Theorem 43 (i)] to the case where $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ may be non-affine maps.
@@
If $\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|\geq 2$, then for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ on $[0,1]$ such that $\displaystyle\sum_{i\in\mathcal{I}_{N}}r_{i}<1+\delta$ and $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$, where we let $r_{i}\coloneqq\|f_{i}\|_{\textup{Lip}}$.
The probability vectors $\mathfrak{p}_{i},i\in\mathcal{I}_{N}$, are determined only by $(\alpha,\beta,\{G_{i}\}_{i},\{H_{i}\}_{i})$ and depend only on $\{g_{i}\}_{i\in\mathcal{I}_{N}}$. The system $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ is not related to this. This contains the case where $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i\in\mathcal{I}_{N}$.
We also deal with the case where $\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$. Let the total variation distance between probability vectors be
$\left\|{\bf p}-{\bf q}\right\|_{1}\coloneqq\sum_{i\in\mathcal{I}_{N}}|p_{i}-q_{i}|,\ \ {\bf p}=(p_{i})_{i},{\bf q}=(q_{i})_{i}\in\overline{\mathcal{P}_{N}}.$
@@
If $\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ on $[0,1]$ such that $\displaystyle\sum_{i\in\mathcal{I}_{N}}r_{i}<1+\delta$, $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$, and $\|{\bf r}-\mathfrak{p}_{0}\|_{1}\geq\epsilon$, where we let $r_{i}\coloneqq\|f_{i}\|_{\textup{Lip}}$ and ${\bf r}\coloneqq(r_{0},\ldots,r_{N-1})$.
We remark that $\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$ if and only if $\mathfrak{p}_{0}=\cdots=\mathfrak{p}_{N-1}$. If $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i\in\mathcal{I}_{N}$, and $\mathfrak{p}_{0}={\bf q}$, then the solution $\varphi$ is smooth on $[0,1]$ and we also have an explicit expression for $\varphi$. See [27, Theorem 43 (ii)].
@@ 3.1. Proofs
We first give notations for symbolic dynamics. Let $\mathbb{N}\coloneqq\{1,2,\ldots\}$. We define a map $\pi\colon(\mathcal{I}_{N})^{\mathbb{N}}\to[0,1]$ by
$\pi(i_{1}i_{2}\ldots)\coloneqq\lim_{k\to\infty}f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0).$
This is well-defined because each $f_{i}$ is a contraction and $\left(f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0)\right)_{k}$ is a Cauchy sequence since
$\left|f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}\circ f_{i_{k+1}}(0)-f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0)\right|\leq\left(\max_{i\in\mathcal{I}_{N}}\|f_{i}\|_{\textup{Lip}}\right)^{k}.$
Let $\xi_{j}\colon(\mathcal{I}_{N})^{\mathbb{N}}\to\mathcal{I}_{N}$ be the projection to the $j$-th coordinate. Let
$I(i_{1},\ldots,i_{n})\coloneqq\left\{w\in(\mathcal{I}_{N})^{\mathbb{N}}\colon\xi_{j}(w)=i_{j},\ 1\leq j\leq n\right\}.$
@@
$\pi\left(I(i_{1},\ldots,i_{k})\right)=\left[f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0),f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(1)\right].$
@@
$\displaystyle\mu_{\varphi}\left(\pi(I(i_{1},\ldots,i_{k}))\right)$ $\displaystyle=\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{k}}(1)\right)-\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{k}}(0)\right)$
$\displaystyle=g_{i_{1}}\circ\cdots\circ g_{i_{k}}(1)-g_{i_{1}}\circ\cdots\circ g_{i_{k}}(0).$
@@
(3.1) $\frac{\nu_{y}(I(i_{1},\ldots,i_{n-1},i_{n}))}{\nu_{y}(I(i_{1},\ldots,i_{n-1}))}=G_{i_{n}}\circ H_{i_{n-1}}\circ\cdots\circ H_{i_{1}}(y),\ n\geq 2.$
Since $\displaystyle\sum_{i\in\mathcal{I}_{N}}G_{i}(y)=1$ for each $y\in Y$, the Kolmogorov consistency condition holds. Therefore, a probability measure on $(\mathcal{I}_{N})^{\mathbb{N}}$ satisfying (3.1) exists uniquely by the Kolmogorov extension theorem [32, Theorem 2.4.3].
For $i\in\mathcal{I}_{N}$, let $F_{i}\colon(\mathcal{I}_{N})^{\mathbb{N}}\to(\mathcal{I}_{N})^{\mathbb{N}}$ be a map defined by $F_{i}(z)=iz$. Now we see that
(3.2) $\nu_{y}=\sum_{i\in\mathcal{I}_{N}}G_{i}(y)\nu_{H_{i}(y)}\circ F_{i}^{-1}.$
Let $\mu_{y}\coloneqq\nu_{y}\circ\pi^{-1}$. Then, $\mu_{y}$ is a Borel probability measure on $[0,1]$. By $\pi\circ F_{i}=f_{i}\circ\pi$ and (3.2),
(3.3) $\mu_{y}=\sum_{i\in\mathcal{I}_{N}}G_{i}(y)\mu_{H_{i}(y)}\circ f_{i}^{-1}.$
@@
Since $H_{i}(0)\neq 0$ can occur, this does not define an iterated function system. The equation (3.3) is in the framework of a place-dependent iterated function system [11].
@@
(3.4) $\nu_{0}\circ\pi^{-1}=\mu_{0}=\mu_{\varphi}.$
Let $\epsilon\in(0,1/2)$. We assume that
(3.5) $\epsilon\leq\min_{i\in\mathcal{I}_{N}}r_{i}\leq\max_{i\in\mathcal{I}_{N}}r_{i}\leq 1-\epsilon.$
Since $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ is a compatible system, $\displaystyle\sum_{i\in\mathcal{I}_{N}}r_{i}\geq 1$ and $r_{i}>0$ for each $i\in\mathcal{I}_{N}$. Let $s\geq 1$ be the constant such that $\displaystyle\sum_{i\in\mathcal{I}_{N}}r_{i}^{s}=1$.
@@
Let $\epsilon_{2}>0$. Then, for $\delta>0$ such that $\displaystyle\sum_{i\in\mathcal{I}_{N}}r_{i}\leq 1+\delta$ and
(3.6) $\frac{\log N}{\log(N/(1+\delta))}<1+\frac{\epsilon_{2}}{\log(1/\epsilon)},$
it holds that
$s\leq\frac{\log N}{\log(N/(1+\delta))}<1+\dfrac{\epsilon_{2}}{\log 1/\epsilon}$
for every choice of $(r_{i})_{i\in\mathcal{I}_{N}}\in\mathcal{P}_{N}$ such that $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$.
@@
Let $\displaystyle u_{i}\coloneqq\frac{r_{i}}{\sum_{i\in\mathcal{I}_{N}}r_{i}}$. Then, $\displaystyle\sum_{i\in\mathcal{I}_{N}}u_{i}=1$. We see that $\displaystyle\sum_{i\in\mathcal{I}_{N}}r_{i}^{s}=1$ if and only if $\displaystyle\sum_{i\in\mathcal{I}_{N}}u_{i}^{s}=\left(\sum_{i}r_{i}\right)^{-s}$. Since $\displaystyle\sum_{i\in\mathcal{I}_{N}}r_{i}\leq 1+\delta$,
$\frac{\log\left(\sum_{i\in\mathcal{I}_{N}}u_{i}^{s}\right)}{s}\geq-\log(1+\delta).$
The sum $\displaystyle\sum_{i\in\mathcal{I}_{N}}u_{i}^{s}$ takes its minimum at $u_{0}=\cdots=u_{N-1}=\dfrac{1}{N}$. We see this by the method of Lagrange multipliers. The minimum of the sum is $N^{1-s}\leq 1$. Hence, $\displaystyle-\frac{s-1}{s}\log N\geq-\log\left(\sum_{i\in\mathcal{I}_{N}}r_{i}\right)$, which is equivalent to $s\leq\dfrac{\log N}{\log(N/(1+\delta))}$. Therefore, if (3.6) holds, then $\displaystyle s\leq\frac{\log N}{\log(N/(1+\delta))}<1+\frac{\epsilon_{2}}{\log(1/\epsilon)}$. ∎
@@
$R_{n}\left({\bf q};z\right)\coloneqq\frac{\nu_{0}\left(I(\xi_{1}(z),\ldots,\xi_{n}(z))\right)}{q_{\xi_{1}(z)}\cdots q_{\xi_{n}(z)}},\ \ {\bf q}=(q_{i})_{i}\in\mathcal{P}_{N},z\in(\mathcal{I}_{N})^{\mathbb{N}}.$
By a measure-theoretic argument, we can show the following assertion:
@@
Let $\widetilde{q}_{i}\coloneqq r_{i}^{s}>0$ and ${\bf\widetilde{q}}\coloneqq(\widetilde{q}_{i})_{i}\in\mathcal{P}_{N}$. If there exists $\epsilon_{2}>0$ such that
(3.7) $\limsup_{n\to\infty}\frac{-\log R_{n}({\bf\widetilde{q}};z)}{n}\leq-\epsilon_{2},\ \textup{ $\displaystyle\nu_{0}$-a.s.$\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
then
@@
We omit the proof of this assertion. See the proof of [27, Lemma 7]. The equation (3.4) plays an important role.
By Lemmas 3.4 and 3.5, we obtain the following assertion:
@@
(3.8) $\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}\leq-\epsilon_{2},\ \textup{ $\displaystyle\nu_{0}$-a.s.$\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
Then, for $\delta=\delta(N,\epsilon,\epsilon_{2})>0$ satisfying (3.6), it holds that $\dim_{H}\mu_{\varphi}<1$.